OpenStudy (anonymous):

Show the following: NOTE: I would like a fair hint please, not the answer. int_{0}^{\infty}x^2e^{-x^2}dx = \frac{1}{2} int_{0}^{\infty}e^{-x^2}dx

OpenStudy (anonymous):

@dpaInc @Mertsj @Zarkon @amistre64 @TuringTest @SmoothMath @lalaly @AccessDenied @aroub @experimentX @eliassaab @FoolForMath

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}x^2e^{-x^2}dx = \frac{1}{2} \int\limits_{0}^{\infty}e^{-x^2}dx\]

OpenStudy (anonymous):

See the second post^

OpenStudy (lalaly):

do u know the gamma function?

OpenStudy (experimentx):

I bet integration by parts first ... then ... then ...

OpenStudy (anonymous):

How? The e^{-x^2} is not integrable (not an elementary function)...

OpenStudy (lalaly):

thats why i asked u if u know the gamma fnction

OpenStudy (experimentx):

It's a definite function ... half normal distribution function has standard value sqrt(pi)/2

OpenStudy (experimentx):

Well ... one way is Gamma function the other way it changing to polar form ... quite a bit messy

OpenStudy (anonymous):

Sorry I do not know gamma function @lalaly. Suppose I know the following: - integration techniques (by parts, U-sub, trig sub) - approximate integration - improper integration - and the pre-calculus stuff Am I able to solve this? I have yet to hear of a gamma function - no gamma functions

OpenStudy (experimentx):

Oh .. sorry, it's *definite integration

OpenStudy (experimentx):

ever done integration in polar coordinates??

OpenStudy (anonymous):

Suppose I do not know that "polar form" you are talking about @experimentX

OpenStudy (anonymous):

Then there is no other way?

OpenStudy (experimentx):

ha ... i meant polar coordinate

OpenStudy (anonymous):

Ok suppose I do not know polar coordinates.

OpenStudy (experimentx):

Oh .. sorry, dxdy = ??

OpenStudy (anonymous):

No.

OpenStudy (experimentx):

\[ dxdy = r drd\theta\]

OpenStudy (anonymous):

No.

OpenStudy (experimentx):

No it's a formula for unit incremental area like dxdy in polar coordinate form

OpenStudy (experimentx):

x = rcos theta y = rsin theta dx dy = r dr d\theta

OpenStudy (experimentx):

you should know these relations

OpenStudy (anonymous):

What is the topic name?

OpenStudy (experimentx):

Ah ... i really don't know .. just hit on google

OpenStudy (anonymous):

Alright leave this question aside for now. I will talk to my instructor about it.

OpenStudy (experimentx):

sure ...